OpenStudy (anonymous):

Thomas has $6.35 in dimes and quarters. The number of dimes is three more than three times the number of quarters. How many quarters does he have? If q represents the number of quarters, then which of the following expressions represents the value of the number of dimes in cents? 3q + 3 6.35 - q 10(3q + 3) Can someone help me work this out please and thanks :)

OpenStudy (anonymous):

For the full solution: 635 = 25q + 10(3q+3) = 55q + 30 605 = 55q q = 11. Thomas has 11 quarters.

OpenStudy (anonymous):

okayy so there is 14d?

OpenStudy (anonymous):

" dimes is three more than three times quarters. " d = 3 + 3 * q

OpenStudy (anonymous):

okayy im sorry i am a little confused i aint good at math

OpenStudy (anonymous):

One of the question is " which of the following expressions represents the value of the number of dimes in cents? "

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

" Thomas has $6.35 in dimes and quarters " 635 = 10d + 25q

OpenStudy (anonymous):

Solve the system of equations